Problem 13-14A Effects of Transactions on Various Ratios [LO13-2]

Denna Company’s working capital accounts at the beginning of the year follow:

   

  Cash $ 71,000    
  Marketable securities $ 22,000    
  Accounts receivable, net $ 352,400    
  Inventory $ 462,600    
  Prepaid expenses $ 8,700    
  Accounts payable $ 201,800
  Notes due within one year $ 102,000    
  Accrued liabilities $ 60,900
During the year, Denna Company completed the following transactions:

   

x. Paid a cash dividend previously declared, $31,000.
a. Issued additional shares of common stock for cash, $202,000.
b. Sold inventory costing $70,800 for $101,000, on account.
c. Wrote off uncollectible accounts in the amount of $10,400, reducing the accounts receivable balance accordingly.
d. Declared a cash dividend, $31,000.
e. Paid accounts payable, $101,600.
f. Borrowed cash on a short-term note with the bank, $61,500.
g. Sold inventory costing $15,420 for $10,280 cash.
h. Purchased inventory on account, $50,750.
i. Paid off all short-term notes due, $163,500.
j. Purchased equipment for cash, $75,800.
k. Sold marketable securities costing $12,000 for cash, $10,000.
l. Collected cash on accounts receivable, $81,100.

   

Required:
1. Compute the following amounts and ratios as of the beginning of the year: (Round your ratios to 2 decimal places.)

         

2.

Indicate the effect of each of the transactions given above on working capital, the current ratio, and the acid-test ratio. Give the effect in terms of increase, decrease, or none. Item (x) is given as an example: Consider each transaction independently and indicate their effects as compared to the ratios and amounts at the beginning of the period.
Answer #1
1(a) Computation of Working Capital
Particulars Amount Amount
Current Assets
Cash $71,000.00
Marketable securities $22,000.00
Accounts receivable, net $352,400.00
Inventory $462,600.00
Prepaid expenses $8,700.00 $916,700.00
Current Liabilities
Accounts payable $201,800.00
Notes due within one year $102,000.00
Accrued liabilities $60,900.00 $364,700.00
Working Capital (A-B) $552,000.00
1(b) Current Ratio= Current Assets/ Current Liabilities
                         = 916700/364700
                         =2.5:1
1( c)Acid test ratio= Liquid Assets / Current Liabilities
=(916700-8700-462600)/364700= 1.22:1
Liquid Assets= Current Assets- Inventory-prepaid expenses
=916700-462600-8700
Part-2 : Statement showing Effect of each Transaction
S.no Effect of each transactions Effect on W.C. Effect on Current ratio Effect on Acid test ratio
1 Paid a cash dividend previously declared, $31,000. No effect Increase Increase
2 Issued additional shares of common stock for cash, $202,000. Increase Increase Increase
3 Sold inventory costing $70,800 for $101,000, on account. Increase Increase Increase
4 Wrote off uncollectible accounts in the amount of $10,400, reducing the accounts receivable balance accordingly. Decrease Decrease Decrease
5 Declared a cash dividend, $31,000. Decrease Decrease Decrease
6 Paid accounts payable, $101,600. No effect Increase Increase
7 Borrowed cash on a short-term note with the bank, $61,500. No effect Decrease Decrease
8 Sold inventory costing $15,420 for $10,280 cash. Decrease Decrease Increase
9 Purchased inventory on account, $50,750. No effect Decrease Decrease
10 Paid off all short-term notes due, $163,500. No effect Increase Increase
11 Purchased equipment for cash, $75,800. Decrease Decrease Decrease
12 Sold marketable securities costing $12,000 for cash, $10,000. Decrease Decrease Decrease
13 Collected cash on accounts receivable, $81,100. No effect no effect no effect